OpenStudy (anonymous):

write the equation of a line in point slope form that passes through the points (6,-8) and (-2,4)

OpenStudy (anonymous):

first yyou need to find the vector of the line. can be found by using: (6, -8) - (-2, 4) = (8 , -12) now you just need to use this expression, FOR EXAMPLE: (x , y) = (6, -8) + k(8 , -12)

OpenStudy (vijay):

2y = -3x + 2

OpenStudy (vijay):

slope is -3/2

OpenStudy (katrinakaif):

First find the slope of the 2 points using (y2-y1)/x2-x1) As (6,-8) (x1,y1) and (-2,4) (x2,y2) Once you get the slope (m) plug into the equation y = mx + b - m bieng the slope then pick one of the coordinates and plug into equation (y-y1) = m (x-x1) and SOLVE =)